Powerball Winner: Take Lump Sum or Take the Annuity So let’s suppose, reader, that you have won Wednesday’s (2016) $1.5 billion Powerball jackpot. Congratulations! You have some important decisions to make, such as what ailing magazine to acquire and what congressional seat your spouse should run for. But first, you must choose whether to take the prize as an annuity paid over 30 years, or a lump-sum payment right now. If I’m reading you right, you should probably take the annuity. First, some background: You might not realize this, but the top prize in the $1.5 billion Powerball is not actually $1.5 billion. If you take the prize as a one-time cash payment, you will get a mere $930 million, before taxes. If you want $1.5 billion, you’ll have to take it in installments over the next 30 years. That’s a long time, and so most people take the cash, according to Kelly Cripe, a Powerball spokeswoman. The Powerball prize is actually what actuaries call an annuity certain: a stream of annual payments, every year from now until 2045, regardless of what happens to you. If you die before 2045, the future payments become part of your estate, like any other asset.

1) What would be a reasonable rate of return to discount the annuity? Explain using financial/economic principles and concepts how you arrive at the rate. Research finance sites on the web (or other places) to document the figures you are using in coming up with the rate.

2) How would you make a decision between these two alternatives: lump-sum or annuity? What would you take into account?

please solve out...
Answer #1

1) Since the cash flows are guaranteed, we can take a small spread over the risk free rate. The risk free rate is 2.5% and if we consider a small spread of 0.5% to reflect some risk in the cash flows, then discount rate is 3.0%

2)

Lump sum is 930 million

PV of 1.5 billion/30 for 30 years is 50 million for 30 yrs

PV = 50+PV(0.03,29,-50)

PV = 50+959.42= 1009.42

So annuity is preferable as it has higher PV